### v1

# abEunwoo Heo    abByeongchan Choi    abJae-Hun Jung Corresponding author

Abstract
Topological data analysis over graph has been actively studied to understand the underlying topological structure of data. However, limited research has been conducted on how different distance definitions impact persistent homology and the corresponding topological inference. To address this, we introduce the concept of path-representable distance in a general form and prove the main theorem for the case of cost-dominated distances. We found that a particular injection exists among the $1$-dimensional persistence barcodes of these distances with a certain condition. We prove that such an injection relation exists for 00- and $1$-dimensional homology. For higher dimensions, we provide the counterexamples that show such a relation does not exist.
Key words.  Topological data analysis; persistent homology; graph data.
Classification codes.  AMS 00A65, 55N31.
1 Introduction
Topological data analysis (TDA), developed recently, has proven its usefulness in various applications [1, 2, 3, 4]. Persistent homology (PH) is one of the primary tools in TDA, using homological characteristics of the data at various scales [2]. The PH method involves constructing the complex in different scales sequentially, capturing how homology changes as the complex is being built and refined. This procedure is known as filtration. This hierarchical changes of homology enables the PH method to provide the topological inference of the data.
For the PH method, it is crucial to transform the given data into a suitable embedding space. Graph representation is widely used in network science for PH analysis [5], including the brain network [6, 7, 8], financial science [9, 10, 11], text mining [12, 13] and music [14, 15].
Once the appropriate transform is identified, applying the PH method requires defining a proper metric for the transformed geometric object, determining the distance between points in the point cloud or between nodes in the graph representation. There is no single way to define this metric or distance as there can be multiple definitions. Definitions of such distances have been empirically established in different fields of study, yet there is a lack of research on how these varying definitions of distance impact PH calculations.
In this paper, we consider a connected weighted graph $G=(V,E,W_{E})$, where $W_{E}:E\rightarrow\mathbb{R}^{>0}$ is a function that assigns weights to the edges. A natural definition of the distance $d_{\text{weight}}(v,w)$ between any two points $v,w\in V$ on the graph can be chosen as the minimum sum of weights along the paths connecting $v$ and $w$. This definition of distance, namely the shortest path distance, satisfies the metric conditions and has been thoroughly studied both in theoretical terms [16, 17, 18] and in computational terms [19, 20, 21], as well as in approximations [22, 23]. The definition is also widely used in broad fields of computer science and is especially powerful in improving the performance of neural networks [24, 25, 26, 27, 28, 29].
On the other hand, the shortest path as the path connecting two vertices with the least number of edges can be also considered. In a similar sense, there are data in which the relevancy between two nodes is more explained in the path passing through the minimum number of edges than in that with minimum sum of weights. In such a case, it is natural to define the distance between two vertices using such paths. The first approach was conducted in previous research [15]. We canonically reformulated the definition of distance $d_{\text{edge}}(v,w)$ between two points on the graph to be the minimum sum of weights between the paths connecting $v$ and $w$ with the minimum number of edges.
In this research we found an interesting relation between $d_{\text{weight}}$ and $d_{\text{edge}}$. That is, for $d_{\text{weight}}\leq d_{\text{edge}}$, we have $\textbf{bcd}_{1}(d_{\text{weight}})\subseteq\textbf{bcd}_{1}(d_{\text{edge}})$, where $\textbf{bcd}_{1}$ is the $1$-dimensional barcode. This property is quite unusual and is generally not expected to hold, as depicted in Figure 1. To explore this, we introduce a more general concept of distance, termed path-representable distance.
Furthermore in this paper, we prove that such an injection relation holds among any path-representable distances for $1$-dimensional homology. This paper is composed of the following sections. In Section 2, explain distinct distance definitions over the given graph, particularly the weight- and edge-minimal distances and the possible relations among them. In Section 3, we introduce the concept of path-representable distance, a generalization of $d_{\text{weight}}$ and $d_{\text{edge}}$ and we introduce the main theory related to path-representable distances. In Section 4, we rigorously prove the main propositions and theorem. In Section 5, we make an experiment to explain how the main theorem works on the higher dimensions. Finally, in Section 6, we provide a brief concluding remark.
2 Different distance definitions and corresponding persistent homology
2.1 Different distance definitions
Let $G=(V,E,W_{E})$ be a connected weighted graph.
Definition 2.1.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Define the distance $d_{\text{weight}}:V\times V\rightarrow\mathbb{R}^{\geq 0}$ such that for any two distinct vertices $v$ and $w$ in $V$,
$d_{\text{weight}}(v,w)=\min_{p}\left\{\sum_{e\in E(p)}W_{E}(e)\mathrel{\bigg{|}}p\text{ is a path from }v\text{ to }w\right\},$
and $d_{\text{weight}}(v,w)=0$ if $v=w$, where $E(p)$ denotes the set of edges composing the path $p$.
Next, consider the definition associated with the paths involved in terms of the minimum number of edges.
Definition 2.2.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Define the distance $d_{\text{edge}}:V\times V\rightarrow\mathbb{R}^{\geq 0}$ such that for any two distinct vertices $v$ and $w$ in $V$,
$d_{\text{edge}}(v,w)=\min_{p}\left\{\sum_{e\in E(p)}W_{E}(e)\mathrel{\bigg{|}}p\text{ is a path between }v\text{ and }w\text{ such that }|E(p)|=m_{vw}\right\},$
and $d_{\text{edge}}(v,w)=0$ if $v=w$, where $E(p)$ denotes the set of edges composing the path $p$, and $m_{vw}$ represents the minimum number of edges among all paths from $v$ to $w$.
Proposition 2.3.
Let $d_{\text{edge}}$ and $d_{\text{weight}}$ be the distances defined above on a connected weighted graph $G=(V,E,W_{E})$. Then the following inequality holds
$d_{\text{weight}}(v,w)\leq d_{\text{edge}}(v,w)\text{ for any }v,w\in V$
Proof.
Take any pair of two distinct vertices $v$ and $w$ in $V$. Let $\mathcal{P}(v,w)$ be the set of all paths between $v$ and $w$. Define $\mathcal{P}_{min}(v,w)$ as the subset of $\mathcal{P}(v,w)$ containing all paths with the minimum number of edges, i.e., $\mathcal{P}_{min}(v,w)=\left\{p\in\mathcal{P}(v,w)\mid|E(p)|=m_{vw}\right\},$ where $m_{vw}$ represents again the minimum number of edges among all paths in $\mathcal{P}(v,w)$. Then, by definition, we have
$d_{\text{weight}}(v,w)=\min_{p\in\mathcal{P}(v,w)}\left\{\sum_{e\in E(p)}W_{E}(e)\right\}\ \leq\min_{p\in\mathcal{P}_{min}(v,w)}\left\{\sum_{e\in E(p)}W_{E}(e)\right\}=d_{\text{edge}}(v,w).$
2.2 Persistent homology
Consider a connected weighted graph $G=(V,E,W_{E})$ with its distance $d:V\times V\rightarrow\mathbb{R}^{\geq 0}$. Define an abstract simplicial complex $\mathbb{X}$ as the power set of the vertex set $V$, denoted by $\mathbb{X}\coloneqq\mathcal{P}(V)$. Consider a real-valued function $f:\mathbb{X}\rightarrow\mathbb{R}$ defined as
$f(\sigma)=\max\{d(v,w)\mid(v,w)\subseteq\sigma\text{ for }v,w\in V\}$
for any $k$-simplex $\sigma\in\mathbb{X}$ whenever $p\geq 1$, otherwise $f(\sigma)=0$. For a sequence of real numbers $0=a_{0}\leq a_{1}\leq\ldots\leq a_{m}$, if we denote $\mathbb{X}_{a_{l}}\coloneqq f^{-1}((-\infty,a_{l}])$, then we have a filtration
$V=\mathbb{X}_{a_{0}}\hookrightarrow\mathbb{X}_{a_{1}}\hookrightarrow\cdots\hookrightarrow\mathbb{X}_{a_{m}}=\mathbb{X}.$
Note that $\mathbb{X}_{a_{0}}=V$ and $\mathbb{X}_{a_{m}}=\mathcal{P}(V)$. This filtration is called the Vietoris-Rips (Rips) filtration. The filtration induced by $f$ allows us to study topological changes in the simplicial complex $\mathbb{X}$ as the parameter $a_{l}$ varies. As the threshold $a_{l}$ increases, more simplices are included in the filtration, revealing more complex interactions among the vertices based on their distances. This process helps to understand the shape of the data represented by the graph $G$.
Persistent homology captures and quantifies these topological changes throughout different scales. For each $k$-dimensional simplex, we track when it appears and becomes part of a larger homological feature as $a_{l}$ increases. This is formally achieved by computing the homology groups $H_{k}(\mathbb{X}_{a_{l}})$ at each stage of the filtration, using homological algebra over the coefficient field $\mathbb{F}$. Persistent homology tracks the birth and death of homological features such as connected components ($k=0$), loops ($k=1$), and voids ($k=2$) as they appear and disappear through the filtration.
A persistence barcode $\textbf{bcd}_{k}(d)$, for each homological dimension $k$, is the visual representations of homological features as a multiset. They can be represented as follows:
$\textbf{bcd}_{k}(d)=\{[a,b]\mid a,b\in\mathbb{R},a<b\}.$
Each barcode consists of intervals, where the beginning of an interval corresponds to a birth (created when a feature appears in the filtration) and the end corresponds to a death (created when a feature is no longer visible in the filtration as a distinct entity). These intervals provide a powerful summary of the topological features of the data across multiple scales, encapsulating both the geometry and topology of the underlying graph structure. By analyzing these barcodes, we can identify significant topological features that persist over a wide range of scales, which correspond to important structural properties of the data. This methodology of tracking and summarizing the change of topological features via persistent homology provides insights into the complex structures hidden in high-dimensional data. In this paper from now on, we fix the coefficient field $\mathbb{F}$ to be $\mathbb{Z}/2\mathbb{Z}$.
In [30], an algorithm for computing persistence barcodes was presented. Let $\mathbf{D}$ be the boundary matrix. Each row and column of the matrix $\mathbf{D}$ correspond to the simplex, and the matrix $(\mathbf{D}_{i,j})$ is defined as $\mathbf{D}_{i,j}=\begin{cases}1&\text{ if }\sigma_{i}\in\partial\sigma_{j}\\ 0&\text{ otherwise}\end{cases}$. The algorithm performs the column operations known as reduction to obtain the reduced binary matrix $\mathbf{R}$, starting from $\mathbf{D}$. Let $\text{low}_{\mathbf{R}}(j)$ represent the row index of the last nonzero element in the $j$th column of $\mathbf{R}$, or be undefined if the $j$th column consists solely of zeros. The column operation algorithm proceeds as follows:
Algorithm 1 $\mathbf{R}=\mathbf{D}\mathbf{V}$ decomposition
[…]
2:  for $j=1$ to $n$ do
3:     while there exists $j^{\prime}<j$ such that $\text{low}_{\mathbf{R}}(j^{\prime})=\text{low}_{\mathbf{R}}(j)$ do
4:        Update $\mathbf{R}[:,j]\coloneqq\mathbf{R}[:,j]-\mathbf{R}[:,j^{\prime}]$
5:     end while
6:  end for
7:  return  $\mathbf{R}$
The resulting matrix $\mathbf{R}$ is called the reduced matrix. Since the algorithm only uses the column operation from left to right, it is equivalent to the decomposition $\mathbf{R}=\mathbf{D}\mathbf{V}$ with an upper triangular matrix $\mathbf{V}$. The important things to know here are that
  1. 1.
    When the $i$ th column of $\mathbf{R}$ corresponding to the $k$-simplex $\sigma_{i}$ is zero, there is a cycle $\tau$ containing $\sigma_{i}$ in $\mathbb{X}_{b}$, where $b$ is the filtration value of $\sigma_{i}$.
  2. 2.
    If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $\sigma_{j}$ causes the cycle $\tau$ to die, where the death time is the value of $\sigma_{j}$. If there is no such $j$, the cycle $\tau$ will persist infinitely.
  3. 3.
    Suppose that the $i$ th column of $\mathbf{R}$ corresponding to the $k$-simplex $\sigma_{i}$ is zero. Let $f$ be the real-valued function induced by the distance $d$ as described in Section 2.2. If there is a $(k+1)$-simplex $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ with $f(\sigma_{i})<f(\sigma_{j})$, then $[f(\sigma_{i}),f(\sigma_{j})]$ is in $\textbf{bcd}_{k}(d)$. Also, if there is no $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$, then $[f(\sigma_{i}),\infty]$ is an element of $\textbf{bcd}_{k}(d)$.
Therefore, the $k$-dimensional persistence barcode $\textbf{bcd}_{k}(d)$ is a multi-set
$\{[f(\sigma_{i}),f(\sigma_{j})]\mid\sigma_{i}\text{ is a $k$-simplex with }i=\text{low}_{\mathbf{R}}(j),f(\sigma_{i})<f(\sigma_{j})\}$
$\bigcup\{[f(\sigma_{i}),\infty]\mid\sigma_{i}\text{ is a $k$-simplex, there isno $\sigma_{j}$ such that }i=\text{low}_{\mathbf{R}}(j)\}.$
When computing the persistence barcode of $\mathbb{X}$, the complexes $\mathbb{X}_{a_{i}}$ are constructed based on the filtration value $a_{i}$. Therefore, simplices with the same value $a_{i}$ appear simultaneously in $\mathbb{X}_{a_{i}}$, yet their order in the column of the boundary matrix $\mathbf{D}$ should differ. However, this order does not affect the computation of the persistence barcode. This property is known as permutation invariance.
Consider the ordering function $\pi$ on $\mathbb{X}$ that determines the order of simplices with the same value for each $\mathbb{X}_{a_{i}}$. We call this function as simplex ordering on $\mathbb{X}$ from the function $f$. Let $\mathbf{D}_{\pi}$ be the boundary matrix with $\pi$. Then $\mathbf{D}_{\pi}$ is uniquely determined if we fix a simplex ordering $\pi$ on $\mathbb{X}$. Also, if $\mathbf{D}_{\pi}$ is uniquely determined, then every $[f(\sigma_{i}),f(\sigma_{j})]$ in $\textbf{bcd}_{k}(d)$ is uniquely determined by $\sigma_{i}$. We obtain the set version of the persistence barcode as
$\textbf{bcd}^{\pi}_{k}(d)=\{[f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}\mid[f(\sigma_{i}),f(\sigma_{j})]\in\textbf{bcd}_{k}(d)\},$
where $\textbf{bcd}^{\pi}_{k}(d)$ denotes the barcode with the ordering function $\pi$. For each element $[f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$ of $\textbf{bcd}^{\pi}_{k}(d)$, $\sigma_{i}$ is called the birth $k$-simplex of $[f(\sigma_{i}),f(\sigma_{j})]_{\sigma_{i}}$.
2.3 Interrelations between two persistence barcodes
Given the connected weighted graph $G=(V,E,W_{E})$, the distances $d_{\text{weight}}$ and $d_{\text{edge}}$ introduced in Section 2.1 are generally distinct. Although $d_{\text{weight}}$ is a metric and $d_{\text{edge}}$ is a semimetric, we found that there is an interesting relation between the $1$-dimensional persistence barcodes by these distances as described in Theorem 2.4.
Theorem 2.4.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Then there exists an injective function $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ defined as $\varphi([a,b])=[a,c]$ such that $b\leq c$ for any $[a,b]\in\textbf{bcd}_{1}(d_{\text{weight}})$.
Figure 1: An example where $d_{1}\leq d_{2}$ yet $\textbf{bcd}_{1}(d_{1})\not\subseteq\textbf{bcd}_{1}(d_{2})$, with $\textbf{bcd}_{1}(d_{1})=\{[4,5]\}$ and $\textbf{bcd}_{1}(d_{2})=\emptyset$. The subgraph $G_{\epsilon}$ of $G$ is defined as $G_{\epsilon}=(V,E_{\epsilon})$, where $E_{\epsilon}=\{e\in E\mid W_{E}(e)\leq\epsilon\}$.
As illustrated in Figure 1, the presence of an inequality between the two arbitrary distances does not guarantee an injection between their $1$-dimensional persistence barcodes in general. However, Theorem 2.4 shows the existence of a pair of distance definitions that could provide an injective relationship in the $1$-dimensional barcode. For example, a possible pair would be $d_{\text{edge}}$ and $d_{\text{weight}}.$
Theorem 2.4 implies that $\textbf{bcd}_{1}(d_{\text{edge}})$ always contains more elements of persistence barcodes than $\textbf{bcd}_{1}(d_{\text{weight}})$. From this perspective, examining $d_{\text{edge}}$ provides access to information that cannot be obtained through $d_{\text{weight}}$ alone. Furthermore, analyzing the significance of the remaining elements of $\textbf{bcd}_{1}(d_{\text{edge}})$ that are not matched with the injective function $\varphi$ in Theorem 2.4 needs further investigation.
Figure 2: An example where $\textbf{bcd}_{1}(d_{\text{edge}})$ and $\textbf{bcd}_{1}(d_{\text{weight}})$ differ yet they form a hierarchical structure within $1$-dimensional hole structures. Bold lines represent newly formed edges, while dashed lines indicate edge that appear at different values.
Figure 2 shows the order in which edges appear in a Rips filtration, based on $d_{\text{edge}}$ and $d_{\text{weight}}$. When calculating $d_{\text{edge}}$, the direct path from $a$ to $b$ yields a value of $5$, but for $d_{\text{weight}}$, the path from $a$ to $b$ via $c$ results in a value of $4$. Due to this difference, $\textbf{bcd}_{1}(d_{\text{edge}})=\{[3,4],[4,5]\}$ and $\textbf{bcd}_{1}(d_{\text{weight}})=\{[3,4]\}$ differ and the injection $\varphi:\textbf{bcd}_{1}(d_{\text{weight}})\rightarrow\textbf{bcd}_{1}(d_{\text{edge}})$ from Theorem 2.4 satisfies $\varphi([3,4])=[3,4]$.
3 Main theory
In this section, we prove Theorem 2.4 by defining more general distance notions and also prove the theorem that is more general than Theorem 2.4. In particular, the distances $d_{\text{weight}}$ and $d_{\text{edge}}$ introduced in Section 2.1 are special cases of a more general distance definition, termed as the path-representable distance, which we will define and explore in this section.
3.1 Path-representable distance
Let $G=(V,E,W_{E})$ be a connected weighted graph. For any two vertices $v$, $w$ in $V$, $P(v,w)$ is the set of paths from $v$ to $w$ that have no repeated vertex, and $\mathcal{P}$ is the union of $P(v,w)$ along all $v,w\in V$. That is,
$\mathcal{P}=\bigcup_{v,w\in V}P(v,w)$
A path choice function is a function $g:V\times V\rightarrow\mathcal{P}$ satisfying $g(a,b)\subseteq g(v,w)\in P(v,w)$ for any vertices $v$ and $w$ in $G$ and any vertices $a$ and $b$ in $g(v,w)$. This property is called consistency.
Definition 3.1.
Consider a connected weighted graph $G=(V,E,W_{E})$. A function $d$ on $V\times V$ is called a path-representable distance if there is a path choice function $g$ such that
$d(v,w)=\sum_{e\in g(v,w)}W_{E}(e)\text{ if }v\neq w$
and $d(v,w)=0$ if $v=w$.
In fact, the consistency property is natural for the definition of the path choice function. The consistency property is equivalent to stating that the chosen path must have an optimal substructure. In computer science, a problem is said to have an optimal substructure if its solution can be constructed by combining solutions to smaller subproblems. For example, in dynamic programming, it is crucial to determine whether or not a given problem possesses an optimal substructure. This characteristic allows a large problem to be divided into smaller sub-problems, each with an optimal solution, which can then be combined to construct the optimal solution for the entire problem, thereby enabling efficient problem resolution.
Consider a connected weighted graph $G=(V,E,W_{E})$ and its path choice function $g$ on $G$. For an edge $e=(v,w)\in E$, we say that $g$ is locally weight-dominated by $e$ if $W_{E}(e)>W_{E}(e^{\prime})$ for any $e^{\prime}\in g(v,w)$. In addition, we say that $g$ is locally cost-dominated by $e$ if $W_{E}(e)\geq\sum_{e^{\prime}\in g(v,w)}W_{E}(e^{\prime})$. Note that if $g$ is locally cost-dominated by $e$, then $g$ is locally weight-dominated by $e$.
We define that $g$ is weight-dominated if $g$ is locally weight-dominated by any $e=(v,w)\in E$ whenever $g(v,w)\neq e$ and $g$ is cost-dominated if $g$ is locally cost-dominated by any $e=(v,w)\in E$ whenever $g(v,w)\neq e$. These conditions mean that there must be a reason to choose a path $g(v,w)$ instead of the single-edge path $e$ in $G$.
Definition 3.2.
(Dominated distance) Consider a connected weighted graph $G=(V,E,W_{E})$ and its path-representable distance $d$ represented by the path choice function $g$ on $G$. We refer to $d$ as weight-dominated if $g$ is weight-dominated and as cost-dominated if $g$ is cost-dominated.
Proposition 3.3.
The two distances $d_{\text{weight}}$ and $d_{\text{edge}}$ defined in Definitions 2.1 and 2.2 are cost-dominated distances.
Proof.
(i) First, we show that $d_{\text{weight}}$ and $d_{\text{edge}}$ are path-representable distances. Consider $d_{\text{weight}}$. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$. Denote $p_{1,k}$ by the minimal path from $v_{1}$ to $v_{k}$ such that $d_{\text{weight}}(v_{1},v_{k})=\sum_{e\in E(p_{1,k})}W_{E}(e)$ for $k=2,3,\ldots,n$. Consider $p_{1,2}$ and $p_{1,3}$. Suppose $p_{1,2}$ and $p_{1,3}$ meet at the midpoint $z\neq v_{1}$. Then, denote the subpath from $v_{1}$ along $p_{1,2}$ to $z$ by $\alpha$, and the subpath from $v_{1}$ along $p_{1,3}$ to $z$ by $\beta$. By the minimality of $p_{1,2}$ and $p_{1,3}$, $\sum_{e\in E(\alpha)}W_{E}(e)=\sum_{e\in E(\beta)}W_{E}(e)$. This is because if one side were smaller, we could find a shorter path by replacing the other. Now, redefine $p_{1,3}$ as a new path starting at $v_{1}$, following $\alpha$ to $z$, and then from $z$ to $v_{3}$ along the path of $p_{1,3}$. Repeat this process until the midpoint where $p_{1,2}$ and $p_{1,3}$ meet disappears.
This process ensures one of the followings: (a) $p_{1,2}$ and $p_{1,3}$ meet only at $v_{1}$ or not at all, or (b) they start from $v_{1}$, overlap only on a common subpath, and do not meet anywhere else until the end. We call the above replacement process a ()(*)( ∗ )-replacement.
Next, perform the ()(*)( ∗ )-replacement for $p_{1,4}$ with respect to $\{p_{1,2},p_{1,3}\}$, and repeat this until $p_{1,n}$ is ()(*)( ∗ )-replaced with respect to $\{p_{1,2},p_{1,3},\ldots,p_{1,n-1}\}$. Then, take $p_{2,1}$, reverse its direction to be the same as $p_{1,2}$, and ()(*)( ∗ )-replace $p_{2,3}$ with respect to $p_{2,1}$. Continue to $p_{2,4}$ with respect to $\{p_{2,1},p_{2,3}\}$, and repeat this process until $p_{2,n}$ is ()(*)( ∗ )-replaced with respect to $\{p_{2,1},p_{2,3},\ldots,p_{2,n-1}\}$. By continuing this process, we ensure that for any distinct $i$ and $j$, the consistency property for $p_{i,j}$ is satisfied. Finally, define the path choice function $g$ as $g(v_{i},v_{j})=p_{i,j}$, thus proving that $d_{\text{weight}}(v_{i},v_{j})=\sum_{e\in g(v_{i},v_{j})}W_{E}(e)$
The ()(*)( ∗ )-replacement process can also be applied in the case of $d_{\text{edge}}$. If there is a midpoint $z$, both the number of edges in the two different paths with the same starting points and the sum of the weights up to the midpoint must be the same. Thus the replacement can be done similarly. Therefore, $d_{\text{edge}}$ becomes the path-representable distance by undergoing the similar procedure.
(ii) Second, we demonstrate that the path choice function $g$ of $d_{\text{weight}}$ in (i) is cost-dominated. Consider any edge $e=(v,w)$ with $g(v,w)\neq e$. Since the edge $e$ itself also constitutes a path from $v$ to $w$, by considering the minimality of $g(v,w)$, we have $W_{E}(e)\geq\sum_{e^{\prime}\in g(v,w)}W_{E}(e^{\prime})$. Additionally, $d_{\text{edge}}$ is cost-dominated by its definition. ∎
3.2 Toy example of path-representable distance
Figure 3: The graph $G=(V,E)$ and all of its possible path choice functions represented by their maximal paths.
Consider a connected weighted graph $G=(V,E,W_{E})$ with $V=\{v_{1},v_{2},v_{3},v_{4}\}$ and $E=$$\{(v_{1},v_{2}),$ $(v_{2},v_{3})$, $(v_{3},v_{4})$,$(v_{1},v_{4})\}$ as shown in the left figure of Figure 3. How many path-representable distances can we make for any path choice function $g$ on $G$? A maximal path can represent any path choice function $g$ on $G$, where the maximal path refers to a path that is maximal with respect to the inclusion. The right figure of Figure 3 represents the maximal paths indicated by the given path choice function on graph $G$. The maximal paths are well-defined due to the consistency property of the path choice function. In this case, a total of eight possible path choice functions can be defined on $G$, as shown in Figure 3.
Figure 4: The weighted graph $G=(V,E,W_{E})$ and all its possible path-representable distances.
For the connected weighted graph $G=(V,E,W_{E})$ shown in the left figure of Figure 4, all path-representable distances derived from all the path choice functions are illustrated in the right figure of Figure 4, with a total of eight possibilities. In this case, all the distances are weight-dominated, while only the distances $d_{5},d_{6},d_{7},d_{8}$ are cost-dominated.
3.3 Main theorem
Similarly, as discussed in Theorem 2.4 with respect to $d_{\text{weight}}$ and $d_{\text{edge}}$, the following generalization can be proved for cost-dominated path-representable distances:
Theorem 3.4.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider two cost-dominated path-representable distances $d_{i}$ and $d_{j}$ by path choice functions $g_{i}$ and $g_{j}$ on $G$, respectively. If $d_{i}(v,w)\leq d_{j}(v,w)$ for any vertices $v$ and $w$, then there exists an injective function $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$ defined as $\varphi_{i,j}([a,b])=[a,c]$ such that $b\leq c$ for any $[a,b]\in\textbf{bcd}_{1}(d_{i})$.
Figure 5: The injective functions $\varphi_{i,j}:\textbf{bcd}_{1}(d_{i})\rightarrow\textbf{bcd}_{1}(d_{j})$, as described in Theorem 3.4, for the four cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ in Figure 4.
Figure 5 illustrates the application of Theorem 3.4 to the cost-dominated path-representable distances $d_{5},d_{6},d_{7},d_{8}$ described in Figure 4. The left figure of Figure 5 shows the partial order relationships among distances $d_{5},d_{6},d_{7},d_{8}$, while the right figure displays the injective functions $\varphi_{5,6},\varphi_{6,7},\varphi_{5,8},\varphi_{8,7}$ between their 1-dimensional persistence barcodes.
3.4 Remarks
The main theorem can also be formulated using the concept of a poset. Let $G$ be a connected weighted graph. To analyze the graph by defining the distance between the nodes, it is natural to define the set whose elements are the distances defined on $G$. Especially in our theory, we focus on the poset $(\mathcal{L},\leq)$ whose elements are the cost-dominated path-representable distances defined on $G$, where for two elements $d_{1},d_{2}\in\mathcal{L}$, $d_{1}\leq d_{2}$ if $d_{1}(v,w)\leq d_{2}(v,w)$ for any two points $v,w\in V$. Furthermore, we calculate persistent homology using the standard algorithm [4], which calculates persistence barcode from the given distance matrix. Note that $(G,d)$ gives a unique distance matrix. We define such a function as $\textbf{bcd}_{k}$, that is, $\textbf{bcd}_{k}(d)$ is a $k$-dimensional persistence barcode obtained by $(G,d)$. Then the codomain of the function will be the poset of persistence barcodes on $G$, as a subset of the poset of interval sets paired with $k$-simplices. Here, the partial order holds if we can find the proper inclusion between the elements with the same associated simplex in the sense of the main theorem. To be precise, for two $k$-dimensional persistence barcode $S_{1}$ and $S_{2}$, we define $S_{1}\leq S_{2}$ if there is an injection $\varphi:S_{1}\hookrightarrow S_{2}$ such that $\varphi([a,b]_{\sigma})=[a,b^{\prime}]_{\sigma}\text{ for any }[a,b]_{\sigma}\in S_{1}\text{ with }b\leq b^{\prime}.$ The main theorem claims that $\textbf{bcd}_{1}$ becomes the order-preserving map.
Note that fixing the vertex order at least once is necessary to identify the exact injection. However, once the injection is identified, it will not change with respect to the change of vertex order. Hence, the whole theory holds when the codomain is a poset of persistence barcodes on $G$ as a subset of the poset of multisets of intervals, carrying the injections defined with the associated birth edge. This gives canonicality to the theory.
Corollary 3.5.
Consider the connected weighted graph $G=(V,E,W_{E})$. Let $\mathcal{L}$ be the poset of path-representable distances from cost-dominated path choice functions on $G$ and $B$ be the poset of persistence barcodes on $G$. There is an order-preserving map $\textbf{bcd}_{1}(\cdot)$ from $\mathcal{L}$ to $B$.
Considering the domain and range, it is natural to question if the least and the greatest elements exist. If we take the domain to be the set of cost-dominated path-representable distances on $G$, then $\textbf{bcd}_{1}$ becomes the order-preserving map. Hence if the least and greatest elements exist, they will map to the least and the greatest elements in the range.
Figure 6: The counterexample for the greatest element in $\mathcal{L}$ illustrated in Proposition 3.6. Here $g_{1}$ and $g_{2}$ are the paths corresponding to $d_{1}$ and $d_{2}$, respectively.
Proposition 3.6.
There is always the least element in the set of cost-dominated path-representable distances on $G$, while there is no greatest element in general.
Proof.
First, if we choose $d_{weight}$ as an element, then it becomes the least element according to the Definition 2.1. Note that $d_{weight}$ is indeed the element in the domain by Proposition 3.3.
Now we take a counterexample for the greatest element. Let $G$ be a cycle graph with at least 5 vertices and suppose that $d_{\max}$ is the greatest element in $P$. As illustrated in Figure 6, fix a control point $v_{0}$. Since $d_{\max}$ is path-representable, for every other vertex $v$ of $G$, we must choose between two paths from $v_{0}$ to $v$ which $d_{\max}$ calculates. Let $v_{i}$ and $v_{i+1}$ be two consecutive nodes. Let $d_{1}$ be the distance that chooses the path from $v_{0}$ to $v_{i}$ that does not contain $v_{i+1}$ and the path from $v_{0}$ to $v_{i+1}$ that contains $v_{i}$. Let $d_{2}$ be the distance that chooses the path from $v_{0}$ to $v_{i+1}$ that does not contain $v_{i}$ and the path from $v_{0}$ to $v_{i}$ that contains $v_{i+1}$. Note that we can make these two distances cost-dominated since $v_{i}$ and $v_{i+1}$ are not adjacent to $v_{0}$, and $G$ is a cycle graph. Then $d_{1}(v_{0},v_{i})>d_{2}(v_{0},v_{i})$ and $d_{1}(v_{0},v_{i+1})<d_{2}(v_{0},v_{i+1})$. Thus, due to the assumption for $d_{\max}$, we have $d_{\max}(v_{0},v_{i})=d_{1}(v_{0},v_{i})$ and $d_{\max}(v_{0},v_{i+1})=d_{2}(v_{0},v_{i+1})$ since there are only two candidates for each selection of path. This leads to a contradiction to the consistency condition of path-representable distance.
Figure 7: $\textbf{bcd}_{1}$ being an order preserving map between the poset of cost-dominated path-representable distance on $G$ (Left) and the poset of $1$-dimensional persistence barcode on $G$ from Figure 4 (Right). Here the arrow denotes the partial order for each poset.
4 Proofs of main theory
Lemma 4.1.
For a weighted graph $G=(V,E,W_{E})$, let $d:V\times V\rightarrow\mathbb{R}$ be any distance defined on $G$. If $\sigma=(v,w)$ be a birth simplex of dimension $1$ with the distance $d(v,w)$ of a $1$-dimensional cycle, there is no vertex $v_{0}\in V(G)$ such that $d(v,v_{0})<d(v,w)$ and $d(w,v_{0})<d(v,w)$.
Proof.
Let $e=(v,w)$ be a birth edge with the distance $d(v,w)$ of a $1$-dimensional cycle. Suppose that there is a vertex $v_{0}\in V(G)$ such that $d(v,v_{0})<d(v,w)$ and $d(w,v_{0})<d(v,w)$. If we denote $t_{0}=\max(d(v,v_{0}),d(w,v_{0}))$, then the edges $(v,v_{0}),(w,v_{0})$ are already in the simplicial complex $\mathbb{X}_{t_{0}}$. Now, considering the $\mathbf{R}=\mathbf{D}\mathbf{V}$ decomposition, with the boundary matrix $D$, the $2$-simplex $(v,w,v_{0})$ maps to $(v,w)+(v,v_{0})+(w,v_{0})$. Of those three edges, $(v,w)$ is the most recent edge born before $(v,w,v_{0})$. Thus $(v,w)$ is necessarily chosen as the index of $\text{low}_{\mathbf{R}}$. The edge $e=(v,w)$ has its birth value as $d(v,w)$ and its death value as $d(v,w)$ because $(v,w)$ and $(v,w,v_{0})$ are created simultaneously. This contradicts the fact that $e$ is a birth edge. ∎
Proposition 4.2.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider any path choice function $g$ on $G$ and its path-representable distances $d$ along $g$. Then every birth simplex of dimension $1$ of $\textbf{bcd}_{1}(d)$ exists in $G$. Therefore, it is justified to call a birth simplex of dimension $1$ as a birth ‘edge’.
Proof.
Let $e=(v,w)$ be a birth simplex of dimension $1$ with a birth time $d(v,w)$. Suppose $e\notin E$. There is a path $g(v,w)$ in $G$ such that $d(v,w)=\sum_{e\in g(v,w)}W_{E}(e)$. Denote the path $g(v,w)$ as the sequence of vertices $v=v_{0},v_{1},v_{2},\ldots,v_{m-1},v_{m}=w$ and $a_{k}=W_{E}((v_{k-1},v_{k}))$. Then, by the consistency property of the path choice function $g$, the path $g(v,w_{m-1})$ is a sequence of $v_{0}$, $v_{1}$, $v_{2}$, $\ldots$, $v_{m-1}$. Therefore, we know that $d(v,v_{m-1})=a_{1}+a_{2}+\cdots+a_{m-1}<d(v,w)$. Since $d(v_{m-1},w)<d(v,w)$ is obvious, this contradicts Lemma 4.1. ∎
Let $\mathcal{B}$ be the set of all birth edges in $\textbf{bcd}_{1}(d)$. By Proposition 4.2, we know that $\mathcal{B}\subseteq E$. Generally, for any edge $e=(v,w)\in E$, it is not necessary that $W_{E}(e)=d(v,w)$. For example, cases where $W(e)>d_{\text{weight}}(v,w)$ are common. However, for the birth edges, the following interesting proposition holds.
Proposition 4.3.
For any $e=(v,w)\in\mathcal{B}$, we always have $W_{E}(e)=d(v,w)$.
Proof.
Consider any birth edge $e=(v,w)\in\mathcal{B}$. Suppose that $W_{E}(e)\neq d(v,w)$. Then the path $g(v,w)$ is not the edge $e$. Write $g(v,w)$ as $v=v_{0}$, $v_{1}$, $\ldots$ , $v_{m}=w$. We know that $d(v,w)=\sum_{e^{\prime}\in g(v,w)}W_{E}(e^{\prime})=\sum_{k=1}^{m}W_{E}((v_{k-1},v_{k}))$. Then $d(v,v_{m-1})=\sum_{k=1}^{m-1}W_{E}((v_{k-1},v_{k}))<d(v,w)$ because the path $g(v_{0},v_{m-1})$ is a sequence $v_{0}$, $v_{1}$, $\ldots$, $v_{m-1}$ and $W_{E}$ is the positive function. Since $d(v_{m-1},w)<d(v,w)$, obviously this contradicts Lemma 4.1. ∎
Consider all path-representable distances on the graph $G=(V,E,W_{E})$. Then, by Propositions 4.2 and 4.3, their birth simplices of dimension $1$ become elements of $E$, and the distances between the endpoints of the edges are given by the weight $W_{E}$.
Let $G=(V,E,W_{E})$ be any weighted graph. A tree of $G$ is a connected subgraph which does not have a cycle. A forest of $G$ is a subgraph which does not have a cycle. A spanning tree of $G$ is a tree of $G$ that includes all the vertices of $G$. For fixed $G$, there are many spanning trees of $G$. A minimum spanning tree of $G$ is a spanning tree of $G$ that has the smallest sum of weights among all possible spanning trees of $G$. Let $\text{MST}(G)$ denote the set of all minimum spanning trees of $G$.
Lemma 4.4.
Let $G=(V,E,W_{E})$ be any connected weighted graph and $\mathcal{T}\in\text{MST}(G)$. Suppose that we have a cycle $\tau$ in $G$ and that there is an edge $e\in\tau$ such that $W_{E}(e)>W_{E}(e^{\prime})$ for any other edge $e^{\prime}\in E$. Then $e\notin\mathcal{T}$.
Proof.
Assume that $e\in\mathcal{T}$. As $\mathcal{T}$ is a tree, if we remove $e$, then $\mathcal{T}$ is split into two subtrees $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. Note that $\tau-e$ is also a path that connects $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. Because $\mathcal{T}$ is a spanning tree, every vertex in $\tau-e$ is in $\mathcal{T}_{1}$ or $\mathcal{T}_{2}$. Therefore, there exists an edge $e^{\prime}$ in $\tau$ connecting $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ such that $W_{E}(e)>W_{E}(e^{\prime})$. However, replacing the edge $e$ in $\mathcal{T}$ with $e^{\prime}$ results in a tree $\mathcal{T}_{1}\cup\mathcal{T}_{2}\cup\{e^{\prime}\}$ that is a spanning tree with a sum of weights smaller than $\mathcal{T}$. This is a contradiction. ∎
Definition 4.5 (Edge swap process of spanning trees).
Let $\mathcal{T}$ be a spanning tree in a connected graph $G$. Consider any edge $e\notin\mathcal{T}$. Then there exists a unique path $p$ in $\mathcal{T}$ connecting the endpoints of $e$. For any edge $e_{p}$ in the path $p$, we can replace the edge $e$ with $e_{p}$. This results in another spanning tree $\mathcal{T}^{\prime}=\mathcal{T}\cup\{e\}-\{e_{p}\}$. This operation is called the edge swap process of a spanning tree.
Consider a connected weighted graph $G=(V,E,W_{E})$. Define $\bar{E}=\{(v,w)\mid v,w\in V\}$. For any path choice function $g$, we have its path-representable distance $d_{g}$ and a weighted complete graph $K^{g}=(V,\bar{E},W_{\bar{E}})$, where $W_{\bar{E}}(e)=d_{g}(v,w)$ for all $e=(v,w)\in\bar{E}$. We name this graph $K^{g}$ the the graph completion of $G$ by $g$.
Proposition 4.6.
Let $G=(V,E,W_{E})$ be a connected weighted graph, and let $K^{g}=(V,\bar{E},W_{\bar{E}})$ be its graph completion by $g$ as explained. If $g$ is weight-dominated, then we have $\text{MST}(K^{g})=\text{MST}(G)$ as the weighted graph.
Proof.
$(\text{MST}(K^{g})\subseteq\text{MST}(G))$. (i) First, we show that any minimum spanning tree of ${K^{g}}$, denoted by $\mathcal{T}$, is a subgraph of $G$ so it is also a spanning tree of $G$. Suppose that there is an edge $e=(v,w)\in\bar{E}$ in $\mathcal{T}\setminus E$. Then $W_{\bar{E}}(e)=d(v,w)=\sum_{e^{\prime}\in g(v,w)}W_{E}(e^{\prime})$, where $g(v,w)$ is different from $e$ because $e$ is not in $G$. Observe that $\tau=g(v,w)+e$ is a cycle. By the consistency of $g$, for any edge $e^{\prime}=(v_{0},w_{0})$ in $g(v,w)$ we know the path $g(v_{0},w_{0})$ is just the edge $(v_{0},w_{0})$, so $W_{\bar{E}}(e^{\prime})=W_{E}(e^{\prime})$. This contradicts Lemma 4.4 because we have $W_{\bar{E}}(e)=\sum_{e^{\prime}\in g(v,w)}W_{E}(e^{\prime})=\sum_{e^{\prime}\in g(v,w)}W_{\bar{E}}(e^{\prime})>W_{\bar{E}}(e^{\prime})$ and $\tau$ is a cycle in $K^{g}$ with $e\in\mathcal{T}$.
(ii) Second, we show that $W_{E}=W_{\bar{E}}$ on $\mathcal{T}$. Consider any edge $e=(v,w)\in\mathcal{T}$. Suppose that $W_{E}(e)\neq W_{\bar{E}}(e)$. This means that there is a path $g(v,w)\neq e$ in $G$. Observe that $\tau=g(v,w)+e$ is a cycle. This contradicts Lemma 4.4 for the same reason as in (i) because $\mathcal{T}$ is a minimum spanning tree of $K^{g}$.
(iii) Consider a minimum spanning tree $\mathcal{T}^{\prime}\in\text{MST}(G)$. If $\mathcal{T}=\mathcal{T}^{\prime}$, then the proposition holds. Suppose that $\mathcal{T}\neq\mathcal{T}^{\prime}$, which means that there is an edge $e\in\mathcal{T}\setminus\mathcal{T}^{\prime}$ and a unique path $p^{\prime}$ in $\mathcal{T}^{\prime}$ connecting the endpoints of $e$. In the path $p^{\prime}$, there is an edge $e^{\prime}\in\mathcal{T}^{\prime}\setminus\mathcal{T}$ because $e+p^{\prime}$ is a cycle in $G$ and $\mathcal{T}$ is a tree. By $(ii)$, we have $W_{E}(e)=W_{\bar{E}}(e)$. Suppose that $W_{E}(e^{\prime})\neq W_{\bar{E}}(e^{\prime})$ and denote $e^{\prime}=(v,w)$. Then there is a path $g(v,w)$ such that $W_{E}(e^{\prime})>W_{E}(e_{i})$ for any $e_{i}\in g(v,w)$. This contradicts Lemma 4.4 because $e^{\prime}\in\mathcal{T}^{\prime}\in\text{MST}(G)$ so that $W_{E}(e^{\prime})=W_{\bar{E}}(e^{\prime})$. Since $\mathcal{T}^{\prime}\in\text{MST}(G)$, we have $W_{E}(e^{\prime})\leq W_{E}(e)$. Otherwise, by replacing $e^{\prime}$ with $e$ in the edge swap process, we have a contradiction as $\mathcal{T}^{\prime}\in\text{MST}(G)$. Similarly, since $\mathcal{T}\in\text{MST}(K^{g})$, we have $W_{\bar{E}}(e^{\prime})\geq W_{\bar{E}}(e)$. To summarize, the followings hold:
  1. 1.
    $W_{E}(e)=W_{\bar{E}}(e)$ and $W_{E}(e^{\prime})=W_{\bar{E}}(e^{\prime})$.
  2. 2.
    $W_{E}(e^{\prime})\leq W_{E}(e)$ and $W_{\bar{E}}(e^{\prime})\geq W_{\bar{E}}(e)$.
Therefore, we conclude that $W_{E}(e)=W_{E}(e^{\prime})$. If we now consider $\mathcal{T}\cup\{e^{\prime}\}-\{e\}$ as a new $\mathcal{T}$, we can repeat the above process. Since this process will terminate after a finite number of steps, we ultimately achieve $\mathcal{T}=\mathcal{T}^{\prime}$, thereby completing the proof.
$(\text{MST}(G)\subseteq\text{MST}(K^{g}))$. The proof can be demonstrated by almost the same method. ∎
Proposition 4.6 implies that any weight-dominated path choice function $g$ always keeps $\text{MST}(G)$ invariant, for which the weight-dominated condition is essential. Otherwise, neither $\text{MST}(G)$ nor $\text{MST}(K^{g})$ can be guaranteed to be larger. A counterexample is shown in Figure 8 below.
Figure 8: A counterexample demonstrating that without the weight-dominated condition, neither $\text{MST}(G)$ nor $\text{MST}(K^{g})$ can be guaranteed to be larger.
In fact, the construction of a minimum spanning tree naturally arises from the standard algorithm discussed in Section 2.2.
Lemma 4.7.
For the fixed order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, consider the set of all $1$-simplices $\sigma_{j}$ such that $i=\text{low}_{\mathbf{R}}(j)$ for some 00-simplex $\sigma_{i}$, denoted by $\mathcal{T}^{\pi}$. Then
$\textnormal{MST}(K^{g})=\{\mathcal{T}^{\pi}\mid\pi\text{ is an order on edge set $\bar{E}$ in $K^{g}$}\}$
Proof.
()superset-of-or-equals($\supseteq$)( ⊇ ) The process of sequentially constructing a minimum spanning tree matches precisely the process of incrementally calculating $\text{low}_{\mathbf{R}}$. To explain in more detail, for every vertex in $K^{g}$, a 00-simplices (vertices) is connected to the $1$-simplices (edges) causing its death. According to the definition of $\text{low}_{\mathbf{R}}$, no cycles can form hence it is a spanning tree. Since $\text{low}_{\mathbf{R}}$ selects edges in order from the lowest value, it becomes the minimum spanning tree. Therefore, for any order $\pi$ on the edge set $\bar{E}$ in $K^{g}$, $\mathcal{T}^{\pi}$ is an element in $\text{MST}(K^{g})$.
()($\subseteq$)( ⊆ ) Consider an arbitrary element $\mathcal{T}\in\text{MST}(K^{g})$. Let $\pi$ be an order on $\bar{E}$ such that, for any given weights on the edges, if the weights are the same, the edges in $\mathcal{T}$ are preferred over those in $K^{g}\setminus\mathcal{T}$. When computing the persistent homology of $K^{g}$, even if the order of the $1$-simplices (edges) corresponding to the indices of the matrix in the $\mathbf{R}=\mathbf{D}\mathbf{V}$ algorithm is rearranged according to $\pi$, the result of the persistent homology computation remains unchanged (permutation invariant). In this case, $\mathcal{T}=\mathcal{T}^{\pi}$. ∎
Proposition 4.8.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider any weight-dominated choice functions $g_{1}$ and $g_{2}$ and theirs completions $K^{g_{1}}$ and $K^{g_{2}}$. Then there exists an order $\pi$ in $\bar{E}$ such that their minimum spanning trees on $K^{g_{1}}$ and $K^{g_{2}}$ induced by the standard algorithm are exactly the same.
Proof.
First, since $g_{1}$ and $g_{2}$ are weight-dominated, by Proposition 4.6, we have $\text{MST}(K^{g_{1}})=\text{MST}(K^{g_{2}})=\text{MST}(G)$. Now, let the minimum spanning trees obtained by the standard algorithm for $K^{g_{1}}$ and $K^{g_{2}}$ be denoted as $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, respectively. Then, by Lemma 4.7, there exists an order $\pi$ on $\bar{E}$ such that $\mathcal{T}_{1}=\mathcal{T}_{2}^{\pi}$. ∎
Proposition 4.9.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider any two path-representable distances $d_{1}$ and $d_{2}$ from two cost-dominated path choice functions $g_{1}$ and $g_{2}$. Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be the set of all birth edges of $\textbf{bcd}_{1}(d_{1})$ and $\textbf{bcd}_{1}(d_{2})$. If $d_{2}(v,w)\leq d_{1}(v,w)$ for any vertices $v$ and $w$, then $\mathcal{B}_{2}\subseteq\mathcal{B}_{1}$ for some simplex ordering $\pi$ on $\mathbb{X}=\mathcal{P}(V)$.
Proof.
For a birth edge $e=\sigma_{i}\in\mathcal{B}_{2}$, suppose that we have a pair of simplices $(\sigma_{i},\sigma_{k})$ such that $\text{low}_{\mathbf{R}_{2}}(k)=i$ for some $\sigma_{k}$, where $\mathbf{R}_{2}=\mathbf{D}_{2}\mathbf{V}_{2}$ in the calculation of $\textbf{bcd}_{1}(d_{2})$. Let $f_{2}$ be a real-valued function induced by $d_{2}$ as in Section 2.2. Since $e$ is a birth edge, $f_{2}(\sigma_{i})<f_{2}(\sigma_{k})$.
(i) First, we show that the $e$-column of the reduced $R_{1}$ is zero. From $e\in\mathcal{B}_{2}$, we know that the $e$-column of $\mathbf{R}_{2}$ is zero. The fact that the $e$-column of $\mathbf{R}_{2}$ is a zero column means that there are column indices $j_{1}$, $\ldots$, $j_{m}$ corresponding to edges $e_{j_{1}},\ldots,e_{j_{m}}$ such that $e+\sum_{l=1}^{m}e_{j_{l}}$ is a cycle and there are 00-simplices $\sigma_{i_{j_{k}}}$ such that $i_{j_{k}}=\text{low}_{\mathbf{R}_{2}}(j_{k})$. By Proposition 4.8, we can assume that the minimum spanning trees $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ for $\mathbf{R}_{1}$ and $\mathbf{R}_{2}$ in the standard algorithm are exactly the same for some fixed order $\pi$ on edge set $\bar{E}$. Therefore, we know that $i_{j_{k}}=\text{low}_{\mathbf{R}_{1}}(j_{k})$ for all $k$. Thus the $e$-column of $\mathbf{R}_{1}$ is zero due to the column indices $j_{1},\ldots,j_{m}$ during the column operations in the standard algorithm.
(ii) Next, we show that if the $1$-cycle $\tau$ corresponding to the birth edge $e=\sigma_{i}$ is terminated by the $2$-simplex $\sigma_{j}$ in $\mathbf{R}_{1}$, then $f_{1}(\sigma_{i})<f_{1}(\sigma_{j})$. Consider a pair of simplices $(\sigma_{i},\sigma_{j})$ which is $\text{low}_{\mathbf{R}_{1}}(j)=i$. For any $t\in\mathbb{R}$, define $\mathbb{X}^{2}_{t}={f_{2}}^{-1}((-\infty,t])$ and $\mathbb{X}^{1}_{t}={f_{1}}^{-1}((-\infty,t])$. As $f_{2}\leq f_{1}$ on $\mathbb{X}$, we know that $\mathbb{X}^{1}_{t}\subseteq\mathbb{X}^{2}_{t}$ for any $t$. In filtration on $\mathbb{X}^{1}$, the $1$-cycle $\tau$ is terminated by $\sigma_{j}$ at a death time $f_{1}(\sigma_{j})$. That is, $\tau$ is zero homologous in $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}$. However, since $\mathbb{X}^{1}_{f_{1}(\sigma_{j})}\subseteq\mathbb{X}^{2}_{f_{1}(\sigma_{j})}$, $\tau$ is also terminated at $f_{1}(\sigma_{j})$ in filtration $\mathbb{X}^{2}$. Therefore, we obtain $f_{2}(\sigma_{k})\leq f_{1}(\sigma_{j})$ because the value $f_{2}(\sigma_{k})$ is a death time of $\tau$, which means the first time to die in filtration $\mathbb{X}^{2}$.
Now, by Proposition 4.3, we know that $W_{E}(e)=W_{\bar{E_{2}}}(e)=f_{2}(e)$ since $e\in\mathcal{B}_{2}$. Then, since $g_{1}$ is cost-dominated, we have $f_{1}(e)=W_{\bar{E}_{1}}(e)\leq W_{E}(e)=f_{2}(e)$. By $d_{2}\leq d_{1}$, we know that $f_{2}(e)\leq f_{1}(e)$ and thus $f_{2}(e)=f_{1}(e)$. Since $e\in\mathcal{B}_{2}$, we have $f_{2}(e)<f_{2}(\sigma_{k})$. Therefore, we obtain $f_{1}(e)=f_{2}(e)<f_{2}(\sigma_{k})\leq f_{1}(\sigma_{j})$.
Finally, by (i) and (ii), an edge $e$ is a birth edge in $\textbf{bcd}_{1}(d_{1})$. This completes the proof. ∎
The following is the proof of the main theorem. See 3.4
Proof.
We show that we can construct an injection from $\textbf{bcd}_{1}(d_{2})$ to $\textbf{bcd}_{1}(d_{1})$. By Proposition 4.8, we can take the order $\pi$ in the edge set $\bar{E}$ so that the minimum spanning trees $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ for $K^{g_{1}}$ and $K^{g_{2}}$ in the standard algorithm are exactly the same. For fixed order $\pi$, we can express any points in $\textbf{bcd}_{1}(d_{2})$ as $[a,b]_{\sigma_{i}}$ corresponding to $(\sigma_{i},\sigma_{k})$ with $i=\text{low}_{\mathbf{R}_{2}}(k)$. Here, $\sigma_{i}$ is a birth simplex of dimension $1$ for the point $[a,b]$. By Proposition 4.9, $\sigma_{i}$ is also a birth simplex of dimension $1$ in $\textbf{bcd}_{1}(d_{1})$, so there is a barcode $[a^{\prime},b^{\prime}]_{\sigma_{i}}$ corresponding to $(\sigma_{i},\sigma_{j})$ such that $i=\text{low}_{\mathbf{R}_{1}}(j)$. We know that $a=a^{\prime}$ because by Proposition 4.3, we have
$a=f_{2}(\sigma_{i})=d_{2}(v,w)=W_{E}(\sigma_{i})=d_{1}(v,w)=f_{1}(\sigma_{i})=a^{\prime},$
where $\sigma_{i}=(v,w)$. Also, in (ii) in the proof of Proposition 4.9, we already show that $b=f_{2}(\sigma_{k})\leq f_{1}(\sigma_{j})=b^{\prime}$. Now, define a function $\varphi$ as $\varphi([a,b]_{\sigma_{i}})=[a,b^{\prime}]_{\sigma_{i}}$ for any $[a,b]\in\textbf{bcd}_{1}(d_{2})$. This map is a well-defined injective function because every point in $\textbf{bcd}_{1}$ is uniquely determined by the birth simplex $\sigma_{i}$ of dimension $1$. This completes the proof. ∎
5 Analysis of dimensional variations of the main theorem
In this section, we briefly discuss whether Theorem 3.4 can be extended to arbitrary $k$-dimensional persistence barcodes. Will it hold for the higher dimensions? Generally, no. Although we were unable to provide a proof for the general case of $k$, we have found counterexamples experimentally demonstrating that the theorem does not hold for $2\leq k\leq 5$.
Statement 5.1.
Let $G=(V,E,W_{E})$ be a connected weighted graph. Consider two cost-dominated path choice functions $g_{1}$ and $g_{2}$ on $G$, and its two path-representable distances $d_{1}$ and $d_{2}$ along $g_{1}$ and $g_{2}$. For any dimension $k$ if $d_{2}(v,w)\leq d_{1}(v,w)$ for any vertices $v,w$, then there exists an injective function $\varphi:\textbf{bcd}_{k}(d_{2})\rightarrow\textbf{bcd}_{k}(d_{1})$.
Table 1: True or False for high dimensional cases $k$ for 5.1
Dimension k 0 1 2 3 4 5
(T)rue or (F)alse T T F F F F
Note: This table shows the results of 5.1 for different dimensional cases. The table indicates whether the theorem holds (True) or does not hold (False) for each dimension.
In the case where $k=0$ in 5.1, $\textbf{bcd}_{k}(d_{\text{edge}})$ and $\textbf{bcd}_{k}(d_{\text{weight}})$ become exactly the same, therefore satisfying the condition. To prove that 5.1 is false for $2\leq k\leq 5$, we performed experiments using Python code. Specifically, knowing that $d_{\text{edge}}\geq d_{\text{weight}}$, we sought to identify the weighted graphs $G=(V,E,W_{E})$ where the condition $\lvert\textbf{bcd}_{k}(d_{\text{edge}})\rvert<\lvert\textbf{bcd}_{k}(d_{\text{weight}})\rvert$ is satisfied for dimension $k$. All codes used in this study are available at https://github.com/AI-hew-math/embedding.
We conjecture that 5.1 is false for any dimension $k\geq 2$ and that a generalized definition of distance with the extended weight, and cost-dominated concepts might exist for higher dimensions, leading to a similar injection relationship as demonstrated for the cases $k=0,1$.
6 Conclusion
This paper addresses unusual relationships that can arise when defining distances on graphs for the computation of persistent homology. Specifically, we explored the distances defined as the minimum sum of weights, $d_{\text{weight}}$, and those prioritizing the minimum number of connecting edges irrespective of their weights, $d_{\text{edge}}$. We discovered that a particular injective function always exists between the 1-dimensional persistence barcodes of these distances. To prove this interesting relationship between such different distances, we introduced the concept of path-representable distance in a general form and extended and proved the main theorem for the case of cost-dominated distances.
The study of how persistent homology changes depending on the chosen distance on a graph can provide insights into how the choice of distance affects the outcomes of homology calculations. Although this paper specifically addresses path-based distances, further research could explore whether similar findings apply to more general distances. These cannot be directly covered by the theory introduced in this paper since the path-representable distance should follow one path and must satisfy the consistency condition. This paper uniquely proves the case for $1$-dimensional homology. We will investigate more general definitions and relationships for higher dimensions in our future research. Finally, the identified injection relations will be examined in applications such as music analysis, which could provide a more practical and insightful interpretative framework.
Acknowledgements: This work was supported by National Research Foundation of Korea under grant number 2018R1D1A1A02086097 and grant number 2021R1A2C3009648. This work was also supported partially by the KIAS Transdisciplinary program grant.